# Task Arithmetic Through The Lens Of One-Shot Federated Learning

### v1

@@ C.1.1 Fine-tuning
C.1.2 Scaling coefficient
To determine the optimal scaling coefficient, we search over the range $[0.05,0.1,0.15,\dots,1.95,2.0]$, selecting the value of $\lambda$ that yields the highest average normalized accuracy on validation datasets.
C.1.3 Hyperparameter for FedGMA
To determine the optimal sign agreement threshold $\rho$ for FedGMA, we search over the range $[0.1,0.2,\dots,1.0]$, selecting the value of $\rho$ that yields the highest average normalized accuracy on validation datasets.
C.1.4 Hyperparameter for CCLIP
To determine the optimal threshold $\rho$ for CCLIP, for each experiment, we search over the range generated by a sequence of five evenly spaced numbers between the minimum task vector norm (including) and maximum task vector norm (excluding) used in the experiment, selecting the value of $\rho$ that yields the highest average normalized accuracy on validation datasets.
C.1.5 Task vector norm
In Table 7, we present the norms of the task vectors utilized in Section 5. Specifically, homogeneous fine-tuning task vectors, as provided by [28], are employed in Section 5.1.2. On the other hand, heterogeneous fine-tuning task vectors, developed as part of our work (detailed in Appendix C.1.1), are used in Section 5.1.1.
@@ C.1.5 Task vector norm
C.2 Additional Experiments Using Task Vectors with Homogeneous Fine-Tuning
In this section, we present additional experiments on task vectors generated through a homogeneous fine-tuning process. These task vectors are provided by [28]. We apply Median, FedGMA, and CCLIP to these task vectors. Due to the homogeneous nature of the fine-tuning process, FedNOVA is not applicable. The hyperparameter search for each method follows the same procedure described in Appendix C.1.
@@ C.2 Additional Experiments Using Task Vectors with Homogeneous Fine-Tuning
Figure 3: Histograms showing the change in average normalized accuracy for three different methods compared to Task Arithmetic by using task vectors with homogeneous fine-tuning. For each plot, the x-axis represents the change in average normalized accuracy, calculated as the difference between the average normalized accuracy of the algorithm used and that of Task Arithmetic. The y-axis indicates the number of experiments within the range of change values. A positive value on the x-axis indicates that the algorithm improves upon Task Arithmetic, while a negative value indicates that the algorithm degrades Task Arithmetic.
Appendix D Merging LLMs
D.1 Additional Experiment Details for Merging LLMs
@@
In order to conduct hyperparameter search, we randomly split $5\%$ of GSM8K, MATH, HumanEval and AlpacaEval into validation datasets. To determine scaling coefficient $\lambda$, we search over the range $[0.2,0.4,0.6,0.8,1.0]$ for Task Arithmetic, FedGMA and CCLIP.
@@
To determine the optimal threshold $\rho$ for CCLIP, for each experiment, we search over the range generated by a sequence of five even spaced numbers between the minimum task vector norm (including) and maximum task vector norm (excluding) used in the experiment.
D.2 Additional Experiment Results on Merging LLMs by Median
@@
Appendix E Generalization Ability of Task Arithmetic
In this section, we explore one potential reason behind the strong empirical performance of Task Arithmetic observed in several experiments especially in scenarios involving LLMs. We conjecture this success is linked to the strong generalization capabilities that Task Arithmetic may inherit from FedAvg, or local SGD.
Research on local SGD has shown that, compared to mini-batch SGD which is the core algorithm used in standard centralized training, local SGD can offer better generalization properties [22, 47, 89]. [47] first observed that switching to local SGD after several epochs of mini-batch SGD training enhances model generalization, leading them to propose a post-local SGD approach. In this scheme, local SGD is only employed in the second training phase, after initial mini-batch SGD training. This two-phase strategy mirrors Task Arithmetic, where we start with a mini-batch pre-trained model, switch to local SGD, and ultimately aggregate the updates.
[22] provided theoretical insights into why local SGD improves generalization. They derived a Stochastic Differential Equation (SDE) that models the long-term behavior of local SGD, observing that it induces a larger drift term compared to standard SGD, thereby adding a regularizing effect. Later on, [89] proved that decentralized SGD is asymptotically equivalent to minimizing the loss function of an average-direction sharpness-aware minimization algorithm, which enhances generalization by seeking flatter regions in the loss landscape. This challenges the common belief that centralized training always outperforms decentralized approaches.
The similar phenomenon has been observed in the context of Task Arithmetic. For example, [86] report in Table 1 that Task Arithmetic occasionally surpasses task-specific fine-tuning. In our experiments, particularly when merging LLMs, Task Arithmetic exhibits strong performance. As shown in Table 5, Task Arithmetic achieves the best results when combining three task vectors. This performance is likely attributed to the remarkable generalization ability of local SGD, even in the one-shot setting of Task Arithmetic.

### v2

@@ B.1.1 Fine-tuning
B.1.2 Scaling coefficient
To determine the optimal scaling coefficient $\lambda$, we searched the range $[0.05,0.1,0.15,\dots,1.95,2.0]$, selecting the value that produces the highest average normalized accuracy in the validation data sets.
B.1.3 Hyperparameter for FedGMA
To determine the optimal sign-agreement threshold $\rho$ for FedGMA, we searched the range $[0.1,0.2,\dots,1.0]$, selecting the value that yields the highest average normalized accuracy in validation datasets.
B.1.4 Hyperparameter for CCLIP
To determine the optimal threshold $\rho$ for CCLIP, for each experiment, we searched the range generated by a sequence of five evenly spaced numbers between the minimum task vector norm (including) and maximum task vector norm (excluding) used in the experiment, selecting the value that yields the highest average normalized accuracy on validation datasets.
B.1.5 Task vector norm
In Table 7, we present the norms of the task vectors used in Section 6. Homogeneous fine-tuning task vectors, as provided by Ilharco et al. (2023), are used in Section 6.1.2, while heterogeneous fine-tuning task vectors, developed as part of our work (detailed in Appendix B.1.1), are used in Section 6.1.1.
@@ B.1.5 Task vector norm
B.2 Additional Experiments Using Task Vectors with Homogeneous Fine-Tuning
In this section, we present additional experiments on task vectors generated through a homogeneous fine-tuning process, provided by Ilharco et al. (2023). We apply Median, FedGMA, and CCLIP to these task vectors. Due to the homogeneous nature of the fine-tuning process, FedNOVA is not applicable. The hyperparameter search for each method follows the same procedure described in Appendix B.1.
@@ B.2 Additional Experiments Using Task Vectors with Homogeneous Fine-Tuning
Figure 3: Histograms showing the change in average normalized accuracy for three different methods compared to Task Arithmetic by using task vectors with homogeneous fine-tuning. The x-axis shows the difference in average normalized accuracy relative to Task Arithmetic, with positive values indicating improvement and negative values indicating degradation. The y-axis represents the number of experiments within each range of change values.
Appendix C Merging LLMs
C.1 Additional Experiment Details for Merging LLMs
@@
In order to conduct hyperparameter search, we randomly split $5\%$ of GSM8K, MATH, HumanEval and AlpacaEval into validation datasets. To determine the optimal scaling coefficient $\lambda$, we searched the range $[0.2,0.4,0.6,0.8,1.0]$ for Task Arithmetic, FedGMA and CCLIP.
@@
To determine the optimal threshold $\rho$ for CCLIP, for each experiment, we searched the range generated by a sequence of five even spaced numbers between the minimum task vector norm (including) and maximum task vector norm (excluding) used in the experiment.
C.2 Additional Experiment Results on Merging LLMs by Median
@@
Appendix D Generalization Ability of Task Arithmetic
In this section, we explore one potential reason behind the strong empirical performance of Task Arithmetic observed in several experiments. We conjecture this success is linked to the strong generalization capabilities that Task Arithmetic may inherit from FedAvg, or local SGD.
Research on local SGD has shown that, compared to mini-batch SGD which is the core algorithm used in standard centralized training, local SGD can offer better generalization properties (Gu et al., 2023; Lin et al., 2018; Zhu et al., 2023). Lin et al. (2018) first observed that switching to local SGD after several epochs of mini-batch SGD training enhances model generalization, leading them to propose a post-local SGD approach. In this scheme, local SGD is only employed in the second training phase, after initial mini-batch SGD training. This two-phase strategy mirrors Task Arithmetic, where we start with a mini-batch pre-trained model, switch to local SGD, and ultimately aggregate the updates.
Gu et al. (2023) provided theoretical insights into why local SGD improves generalization. They derived a Stochastic Differential Equation (SDE) that models the long-term behavior of local SGD, observing that it induces a larger drift term compared to standard SGD, thereby adding a regularizing effect. Later on, Zhu et al. (2023) proved that decentralized SGD is asymptotically equivalent to minimizing the loss function of an average-direction sharpness-aware minimization algorithm, which enhances generalization by seeking flatter regions in the loss landscape. This challenges the common belief that centralized training always outperforms decentralized approaches.
The similar phenomenon has been observed in the context of Task Arithmetic. For example, Yu et al. (2024) report in Table 1 that Task Arithmetic occasionally surpasses task-specific fine-tuning. In our experiments, particularly when merging LLMs, Task Arithmetic exhibits strong performance. As shown in Table 5, Task Arithmetic achieves the best results when combining three task vectors. This performance is likely attributed to the remarkable generalization ability of local SGD, even in the one-shot setting of Task Arithmetic.